# Measuring the Temperature of Extremely Hot Shock-Heated Gas in the Major Merger MACS J0717.5++3745 With Relativistic Corrections to the Sunyaev-Zel’dovich Effect

### v1

@@ 3.3 Unwanted Astrophysical Emission
3.3.1 Source Catalog Generation
@@
PCAT-DE uses a Bayesian hierarchical model to jointly fit the flux density and position of point-like sources, the fluctuation and amplitude of diffuse foreground/background components, and the amplitude of surface brightness templates over a set of multi-band photometric maps (Feder et al., 2023). In this analysis, we use PCAT-DE to recover the point source SEDs from the Herschel-SPIRE photometer maps of MACS J0717.5++3745.
@@
To construct a point source catalog, MACS J0717.5++3745 photometer images in all three Herschel-SPIRE bands (600, 856, and 1200 GHz) are modeled using a modified version of the procedure developed for Butler et al. (2022). Because our goal is to recover a model of the sources comprising the CIB, we fix the SZe template amplitudes for the PMW and PLW bands to their fiducial values from Bolocam measurements of the optical depth (Sayers et al., 2019) and the X-ray inferred temperature from XMM-Newton2.4) so that the SZe is no longer a free parameter. Point sources, the cirrus foreground, and an overall offset are then fit over a $10^{\prime}\times 10^{\prime}$ region centered on the X-ray centroid of the cluster, as described in Feder et al. (2023). This results in a PCAT source catalog describing the positions and fluxes of point sources that is decoupled from diffuse sources of emission in the cluster field.
@@
Residuals between the input Herschel-SPIRE photometer maps and the CIB point source model constructed by PCAT are shown in Figure 7. The CIB model removes point source emission from all three bands and leaves a strong diffuse signal that is brightest in the $600$ GHz band, consistent with the SZe. A false three-color image of the residuals that can be directly compared to the upper left panel of Figure 2 is also shown in Fig. 7. This visually illustrates our high-SNR measurement of diffuse SZe emission in MACS J0717.5++3745 with the Herschel-SPIRE photometer.
@@
Figure 7: Herschel-SPIRE photometric imaging of MACS J0717.5++3745. From top to bottom the rows show PSW ($1200$ GHz), PMW ($856$ GHz), and PLW ($600$ GHz). From left to right the columns show the observed image centered on MACS J0717.5++3745, a fitted CIB point source model constructed using PCAT-DE (Feder et al., 2020), and the residual after subtraction of this model. The pointing of each of the seven detectors are shown as black dots. This demonstrates that we are able to accurately remove CIB contamination while preserving the SZe signal, as primarily seen in the $600$ GHz panel. The bottom panel is a false color image of the three residuals illustrating our high-SNR measurement of a diffuse signal consistent with the extremely red spectrum expected from the SZe.
3.3.2 CIB Correction
@@
With a reliable point source catalog, we can next model the spectrum of each point source. A greybody spectrum (e.g., Planck Collaboration et al., 2014) is fit to the SED of each photometer source identified by PCAT,
@@
where the free parameter $\epsilon$ describes the amplitude of the source in mJy and $T_{dust}$ the temperature of the dust. We set the pivot frequency at $\nu_{p}=856$ GHz and fix $\beta=1.65$ to match the Planck model (Planck Collaboration et al., 2016). The fits are implemented via a Monte Carlo markov chain (MCMC) sampler (Foreman-Mackey et al., 2013) with loss function given by the $\chi^{2}$ distribution. Uniform priors of $T_{dust}\in[5,200]$ K and $\epsilon\in[0,10^{5}]$ mJy ensure that these fits remain within a physically meaningful space of values. The fit outputs are posterior distributions for $\epsilon$ and $T$ for each of the PCAT-detected sources in the cluster region. On a source-by-source basis SEDs from these posterior predictive distributions are sampled in the SPIRE-FTS spectral binning, which allows us to create a probability distribution of flux densities for each source $\{F_{\nu,s}\}$.
@@
From these probability densities, we generate mock catalogs and project them into corresponding mock images. We create a set of 5,000 such mock catalogs to better sample the source to source covariance derived from the PCAT catalog posterior. For each catalog, we assume that each source is point-like and convolve with the frequency dependent SPIRE-FTS PSF to obtain a mock map. Each source’s location is fixed to the median position obtained from PCAT-DE. We then extract the observed flux density in each of the $28$ SPIRE-FTS pointing sets for jiggle position 0. The flux density from CIB sources is taken as the ensemble median over the set of 5,000 images in each pointing. These spectra are then averaged per detector, and the uncertainties of the averaged pointings are added in quadrature, the CIB estimates for each detector are shown in Figure 8. For each CIB model a diagonal covariance matrix is formed with diagonal elements given by the uncertainty estimated above. After subtraction of the CIB model, the final noise covariance matrix for these observations is formed from the sum of the noise covariance matrix estimated from the jackknives as described in Section 3.2 and the posterior variance of the CIB model.
@@
Figure 8: Residuals of the measured spectra and CIB model for each of the seven detectors used in this analysis. The purple points show the Dark Sky template-subtracted signal shown in Fig. 4, the blue curve is the estimated CIB spectrum with width set by the 68% model uncertainty, and the green points are the CIB-subtracted measurements. The CIB corrections act to preferentially reduce flux at high frequencies, as one would expect from the typical greybody model of a dusty infrared galaxy at this redshift.
4 SZe Modeling and Inference
Figure 9: Best-fitting curves to the SZe spectrum from the combined SPIRE-FTS and Bolocam intensity measurements (pink points). The median fit is shown in black, with $1\sigma$ confidence intervals in blue and $2\sigma$ confidence intervals in light blue. The black data points show the SPIRE-FTS data averaged over a correlation length of six spectral channels. In the right hand panel we show the difference between the median fit and the data normalized by the uncertainty at each point.
@@ 4 SZe Modeling and Inference
Table 3: Maximum a posteriori (MAP) values for the parameters of our fit to each of the seven SPIRE-FTS detectors. Error bars indicate 68 per cent credible regions.
@@
where $v_{\text{pec}}$ is the peculiar velocity of the ICM, $y$ is the total Comptonization parameter, and $T_{\rm rSZe}$ is the ICM temperature estimated by the rSZe, which is assumed to be constant along the line of sight. To marginalize over the absolute signal baseline of the SPIRE-FTS data, which is unconstrained, we include the term $O_{FTS}$ that describes an additive offset to the sub-mm data. We jointly fit the CIB-corrected SPIRE-FTS spectra shown in Figure 8 and the Bolocam $140$ and $270$ GHz data (described in §2.2) to this model to constrain the free parameters. The nuisance parameter $O_{FTS}$ is fit to the SPIRE-FTS data points only, and does not contain information of astrophysical importance. The data is fit with a MCMC sampler that evaluates $\Delta I_{kSZe}$ and $\Delta I_{tSZe}$ for each parameter combination using SZpack (Chluba et al., 2012, 2013). As part of this fit, we include uniform priors on $T_{\rm rSZe}$$\in[0,60]$ keV and $y\in[0,1]$. The lower bounds are designed to ensure the MCMC does not sample negative values, which are not physically allowed. The upper bounds are designed to be minimally informative, and do not noticeably impact the recovered posteriors. We also include an informative Gaussian prior on the value of $v_{\text{pec}}$ based on the cluster-member redshifts derived in §2.3 and quantified in Table 2, as well as absolute bounds of $v_{\text{pec}}$ c-1 $\in[-0.1,0.1]$.
@@
The MCMC chains are initialized using random draws of the parameter values that sample the parameter space near the values of $T_{\rm rSZe}$$v_{\text{pec}}$  and $y$ predicted from the fiducial values in Table 2. The initial draws for the nuisance parameter $O_{FTS}$ are $\in[-2.5,2.5]$ $\text{MJySr}^{-1}$. We run each of these chains for twenty thousand steps, and we cut the first three thousand steps as the burn-in samples. A Gelman-Rubin statistic (Gelman & Rubin, 1992) is used to assess the convergence of the chains, where we find that $\hat{R}=1.0$ and only has variation across all detector fits at the level of $10\times 10^{-5}$. The marginalized posterior distributions of these fits are shown in Figures 9 and 10, and the maximum a posteriori (MAP) parameter values for each SPIRE-FTS detector are listed in Table 3.
@@
To assess potential biases in our fitting procedure, we created 100 mock data sets of the rSZe data with realistic noise generated through a Cholesky decomposition. For each mock, the input value of $T_{\rm rSZe}$ is drawn from a Gaussian distribution with $\mu=19.5$ keV and $\sigma=12$ keV. The other parameters are drawn from the posterior distributions shown in Figure 10 and in Table 3. Each of these mocks is then to fit in the same manner as the SPIRE-FTS data, producing a set of $100$ best fit temperatures. We re-sample these outputs with a jackknife method, wherein $10^{4}$ sets of seven measurements are aggregated via the LRGS software package, following the same methodology as applied to the observed data. For each aggregate posterior we extract the MAP value and produce a set of $10^{4}$ estimates. These values are then compared to the mean temperature of $19.5$ keV, and we find an on-average multiplicative bias of $1.016\begin{subarray}{c}+0.003\\ -0.003\end{subarray}$. We conclude that the bias in $T_{\rm rSZe}$ from our fitting methodology is negligible compared to measurement uncertainties. Further, the recovered uncertainties on $T_{\rm rSZe}$ in the mock estimates are comparable to that of the recovered data.
5 Discussion
@@
Data from the SPIRE-FTS in combination with ground-based SZe data from Bolocam have permitted a high significance measurement of the ICM temperature in the quadruple merger MACS J0717.5++3745. We find a cluster-average temperature of $T_{\rm rSZe}$$15.1\begin{subarray}{c}+3.8\\ -3.3\end{subarray}$ keV with an intrinsic scatter over the cluster face corresponding to $\sigma_{\rm rSZe}$ $5.4\begin{subarray}{c}+5.1\\ -3.4\end{subarray}$ keV. This work better establishes the rSZe as a viable tool for probing ICM thermodynamics, particularly of the extremely hot gas expected in major mergers (and confirmed in MACS J0717.5++3745 based on our analysis). In addition, while MACS J0717.5++3745 is a moderate redshift source, our analysis also serves to advance the rSZe as a potential path to probing the ICM in such systems at higher redshift.
@@
It is informative to compare the ICM temperature we obtain from the rSZe against temperatures obtained from X-ray facilities. Differences between them may be expected from at least two sources. First, the differing line-of-sight weighting of the signals combined with non-isothermal distributions within the ICM can lead to discrepancies. Second, the presence of measurement systematics, in particular X-ray calibration uncertainties and band limitations, can also lead to different inferences about ICM temperature.
@@
Regarding the first possibility, the rSZe measurement is effectively pressure-weighted so it is biased toward any gas in post-shock regions that has extreme temperatures and pressures (e.g., Mroczkowski et al., 2019). In contrast, the X-ray signal is biased toward high-density regions, which often correspond to low-temperature remnant cool cores in merging systems (e.g., Rossetti & Molendi, 2010). In addition, differences are possible even in more quiescent systems, primarily as a result of cooling near the cluster center resulting in an enhanced gas density within that region (e.g., Hudson et al., 2010). As a consequence, the X-ray signal will be biased toward these cooler central regions, which would result in a lower temperature compared to that derived from the rSZe.
@@
Regarding the second possibility, we start by noting that there is a well established temperature calibration discrepancy between XMM-Newton and Chandra, which was studied in particular detail by Schellenberger et al. (2015), who established that at least one satellite has a large systematic error in its calibration. In general, they find that Chandra obtains higher temperatures than XMM-Newton, with a fractional difference of approximately 5–20 per cent for typical ICM temperatures, with larger fractional differences at higher temperatures. This difference provides a characteristic scale for the possible systematic error that may exist in their temperature measurements. While more difficult to quantify, we also reiterate the band limitations of the X-ray facilities noted in Section 1, which render them largely insensitive to photon energies $\gtrsim 8$ keV. Thus, both XMM-Newton and Chandra have little to no sensitivity to the hottest regions that may be present in the ICM, which could produce lower X-ray-derived temperatures for such regions.

### v2

@@ III.3 Unwanted Astrophysical Emission
III.3.1 Source Catalog Generation
@@
PCAT-DE uses a Bayesian hierarchical model to jointly fit the flux density and position of pointlike sources, the fluctuation and amplitude of diffuse foreground/background components, and the amplitude of surface brightness templates over a set of multiband photometric maps (Feder et al., 2023). In this analysis, we use PCAT-DE to recover the point source SEDs from the Herschel-SPIRE photometer maps of MACS J0717.5++3745.
@@
To construct a point source catalog, MACS J0717.5++3745 photometer images in all three Herschel-SPIRE bands (600, 856, and 1200 GHz) are modeled using a modified version of the procedure developed for Butler et al. (2022). Because our goal is to recover a model of the sources comprising the CIB, we fix the SZe template amplitudes for the photometer medium wavelength (PMW) and photometer long wavelength (PLW) bands to their fiducial values from Bolocam measurements of the optical depth (Sayers et al., 2019) and the X-ray-inferred temperature from XMM-NewtonII.5) so that the SZe is no longer a free parameter. Point sources, the cirrus foreground, and an overall offset are then fit over a $10^{\prime}\times 10^{\prime}$ region centered on the X-ray centroid of the cluster, as described in Feder et al. (2023). This results in a PCAT source catalog describing the positions and fluxes of point sources that is decoupled from diffuse sources of emission in the cluster field.
@@
Residuals between the input Herschel-SPIRE photometer maps and the CIB point source model constructed by PCAT are shown in Figure 7. The CIB model removes point source emission from all three bands and leaves a strong diffuse signal that is brightest in the $600$ GHz band, consistent with the SZe. A false three-color image of the residuals that can be directly compared to the upper left panel of Figure 2 is also shown in Fig. 7. This visually illustrates our high-S/N measurement of diffuse SZe emission in MACS J0717.5++3745 with the Herschel-SPIRE photometer.
@@
Figure 7: Herschel-SPIRE photometric imaging of MACS J0717.5++3745. From top to bottom, the rows show PSW ($1200$ GHz), PMW ($856$ GHz), and PLW ($600$ GHz). From left to right, the columns show the observed image centered on MACS J0717.5++3745, a fitted CIB point source model constructed using PCAT-DE (Feder et al., 2020), and the residual after subtraction of this model. The pointing of each of the seven detectors is shown with their respective symbols as indicated in Figures 2,4,8, and 9. This demonstrates that we are able to accurately remove CIB contamination while preserving the SZe signal, as primarily seen in the $600$ GHz panel. The bottom panel is a false color image of the three residuals with contours overlaid from the Bolocam $140$ GHz measurements. This illustrates our high-S/N measurement of a diffuse signal that is consistent with the expected red spectrum of the SZe. .
III.3.2 CIB Correction
@@
With a reliable point source catalog, we can next model the spectrum of each point source. A graybody spectrum (e.g., Planck Collaboration et al., 2014) is fit to the SED of each photometer source identified by PCAT,
@@
where the free parameter $\epsilon$ describes the amplitude of the source in mJy and $T_{dust}$ the temperature of the dust. We set the pivot frequency at $\nu_{p}=856$ GHz and fix $\beta=1.65$ to match the Planck model (Planck Collaboration et al., 2016). The fits are implemented via a Monte Carlo Markov Chain (MCMC) sampler (Foreman-Mackey et al., 2013) with loss function given by the $\chi^{2}$ distribution. Uniform priors of $T_{dust}\in[5,200]$ K and $\epsilon\in[0,10^{5}]$ mJy ensure that these fits remain within a physically meaningful space of values. The fit outputs are posterior distributions for $\epsilon$ and $T$ for each of the PCAT-detected sources in the cluster region. On a source-by-source basis, SEDs from these posterior predictive distributions are sampled in the SPIRE-FTS spectral binning, which allows us to create a probability distribution of flux densities for each source $\{F_{\nu,s}\}$.
@@
From these probability densities, we generate mock catalogs and project them into corresponding mock images. We create a set of 5,000 such mock catalogs to better sample the source-to-source covariance derived from the PCAT catalog posterior. For each catalog, we assume that each source is pointlike and convolve with the frequency-dependent SPIRE-FTS PSF to obtain a mock map. Each source’s location is fixed to the median position obtained from PCAT-DE. We then extract the observed flux density in each of the $28$ SPIRE-FTS pointing sets for jiggle position 0. The flux density from CIB sources is taken as the ensemble median over the set of $5000$ images in each pointing. These spectra are then averaged per detector, and the uncertainties of the averaged pointings are added in quadrature. The CIB estimates for each detector are shown in Figure 8. For each CIB model, a diagonal covariance matrix is formed with diagonal elements given by the uncertainty estimated above. After subtraction of the CIB model, the final noise covariance matrix for these observations is formed from the sum of the noise covariance matrix estimated from the jackknives as described in Section III.2 and the posterior variance of the CIB model.
@@
Figure 8: Residuals of the measured spectra and CIB model for each of the seven detectors used in this analysis. The purple points show the Dark Sky template-subtracted signal shown in Fig. 4, the blue curve is the estimated CIB spectrum with width set by the 68% model uncertainty, and the green points are the CIB-subtracted measurements. All panels share the same $x$ and $y$ axes, which are depicted only for the lower-left subplot. The CIB corrections act to preferentially reduce flux at high frequencies, as one would expect from the typical graybody model of a dusty infrared galaxy at this redshift.
IV SZe Modeling and Inference
Figure 9: Shown here are best-fitting curves to the SZe spectrum from the combined SPIRE-FTS (right) and Bolocam intensity measurements (left). The median fit is shown in black, with a $2\sigma$ confidence interval in green. The black data points on the right plots show the average of the SPIRE-FTS data, taken over a correlation length of six spectral channels. The uncertainty on these points is estimated by $\left(I^{T}CI\right)^{-1/2}$, where $I$ is the identity matrix, and $C$ is the covariance matrix sliced into $6\times 6$ cutouts. The highest-frequency point is averaged over five spectral channels. The reduced $\chi^{2}$ values computed with the full covariance matrix are shown in each panel and in Table 3. They imply much better fits than would be inferred directly from the pink data points due to the large correlations between adjacent bins. All panels share the same $x$ and $y$ axes which is depicted only for the lower-left subplots.
@@ IV SZe Modeling and Inference
$\chi^{2}$ 0.920 1.421 1.245 1.785 1.740 1.589 2.537
Table 3: Maximum a posteriori values for the parameters of our fit to each of the seven SPIRE-FTS detectors. Error bars indicate 6868% credible regions.
@@
where $v_{\text{pec}}$ is the peculiar velocity of the ICM, $y$ is the total Comptonization parameter, and $T_{\rm rSZe}$ is the ICM temperature estimated by the rSZe, which is assumed to be constant along the line of sight. To marginalize over the absolute signal baseline of the SPIRE-FTS data, which is unconstrained, we include the term $O_{FTS}$ that describes an additive offset to the submillimeter data. We jointly fit the CIB-corrected SPIRE-FTS spectra shown in Figure 8 and the Bolocam $140$ and $270$ GHz data (described in §II.3) to this model to constrain the free parameters. The nuisance parameter $O_{FTS}$ is fit to the SPIRE-FTS data points only, and does not contain information of astrophysical importance. The data is fit with a MCMC sampler that evaluates $\Delta I_{kSZe}$ and $\Delta I_{tSZe}$ for each parameter combination using SZpack (Chluba et al., 2012, 2013). As part of this fit, we include uniform priors on $T_{\rm rSZe}$$\in[0,60]$ keV and $y\in[0,1]$. The lower bounds are designed to ensure the MCMC does not sample negative values, which are not physically allowed. The upper bounds are designed to be minimally informative and do not noticeably impact the recovered posteriors. We also include an informative Gaussian prior on the value of $v_{\text{pec}}$ based on the cluster-member redshifts derived in §II.4 and quantified in Table 2, as well as absolute bounds of $v_{\text{pec}}$ c-1 $\in[-0.1,0.1]$. The way that HIPE processes the FTS data produces spectra that have an additive signal offset that is almost entirely determined by time-dependent properties of the instrument. While it is possible to place a prior on this offset based on expected total signal from the SZe, CIB, cirrus, and other relevant astrophysical sources, we found that these signal levels do not describe the observed offsets, so such a prior is not informative for our analysis. We thus treat the offset as a free parameter.
@@
The MCMC chains are initialized using random draws of the parameter values that sample the parameter space near the values of $T_{\rm rSZe}$, $v_{\text{pec}}$, and $y$ predicted from the fiducial values in Table 2. The initial draws for the nuisance parameter $O_{FTS}$ are $\in[-2.5,2.5]$ $\text{MJySr}^{-1}$. We run each of these chains for twenty thousand steps, and we cut the first three thousand steps as the burn-in samples. A Gelman-Rubin statistic (Gelman and Rubin, 1992) is used to assess the convergence of the chains, where we find that $\hat{R}=1.0$ and only has variation across all detector fits at the level of $10\times 10^{-5}$. The posterior predictive from these fits is shown in Figure 9, and the posterior constraints are shown in Figure 10 and in Table 3.
@@
To assess potential biases in our fitting procedure, we created 100 mock data sets of the rSZe data with realistic noise generated through a Cholesky decomposition. For each mock, the input value of $T_{\rm rSZe}$ is drawn from a Gaussian distribution with $\mu=19.5$ keV and $\sigma=12$ keV. The other parameters are drawn from the posterior distributions shown in Figure 10 and in Table 3. Each of these mocks is then to fit in the same manner as the SPIRE-FTS data, producing a set of $100$ best-fit temperatures. We resample these outputs with a jackknife method, wherein $10^{4}$ sets of seven measurements are aggregated via the LRGS software package, following the same methodology as applied to the observed data. For each aggregate posterior, we extract the MAP value and produce a set of $10^{4}$ estimates. These values are then compared to the mean temperature of $19.5$ keV, and we find an on-average multiplicative bias of $1.016\begin{subarray}{c}+0.003\\ -0.003\end{subarray}$. We conclude that the bias in $T_{\rm rSZe}$ from our fitting methodology is negligible compared to measurement uncertainties. Further, the recovered uncertainties on $T_{\rm rSZe}$ in the mock estimates are comparable to those of the recovered data.
V Discussion
@@
Data from the SPIRE-FTS in combination with ground-based SZe data from Bolocam have permitted a high-significance measurement of the ICM temperature in the quadruple merger MACS J0717.5++3745. We find a cluster-average temperature of $T_{\rm rSZe}$$15.1\begin{subarray}{c}+3.8\\ -3.3\end{subarray}$ keV with an intrinsic scatter over the cluster face corresponding to $\sigma_{\rm rSZe}$ $5.4\begin{subarray}{c}+5.1\\ -3.4\end{subarray}$ keV. This work better establishes the rSZe as a viable tool for probing ICM thermodynamics, particularly of the extremely hot gas expected in major mergers (and confirmed in MACS J0717.5++3745 based on our analysis). In addition, while MACS J0717.5++3745 is a moderate redshift source, our analysis also serves to advance the rSZe as a potential path to probing the ICM in such systems at higher redshift.
V.1 Potential Differences between rSZE- and X-ray-Measured ICM Temperatures
@@
It is informative to compare the ICM temperature we obtain from the rSZe against temperatures obtained from X-ray facilities. Differences between them may be expected from at least two sources. First, the differing line-of-sight weighting of the signals combined with nonisothermal distributions within the ICM can lead to discrepancies. Second, the presence of measurement systematics, in particular X-ray calibration uncertainties and band limitations, can also lead to different inferences about ICM temperature.
@@
Regarding the first possibility, the rSZe measurement is effectively pressure-weighted so it is biased toward any gas in post-shock regions that has extreme temperatures and pressures (e.g., Mroczkowski et al., 2019). In contrast, the X-ray signal is biased toward high-density regions, which often correspond to low-temperature remnant cool cores in merging systems (e.g., Rossetti and Molendi, 2010). In addition, differences are possible even in more quiescent systems, primarily as a result of cooling near the cluster center, resulting in an enhanced gas density within that region (e.g., Hudson et al., 2010). As a consequence, the X-ray signal will be biased toward these cooler central regions, which would result in a lower temperature compared to that derived from the rSZe.
@@
Regarding the second possibility, we start by noting that there is a well-established temperature calibration discrepancy between XMM-Newton and Chandra, which was studied in particular detail by Schellenberger et al. (2015), who established that at least one satellite has a large systematic error in its calibration. In general, they find that Chandra obtains higher temperatures than XMM-Newton, with a fractional difference of approximately 552020% for typical ICM temperatures, with larger fractional differences at higher temperatures. This difference provides a characteristic scale for the possible systematic error that may exist in their temperature measurements. While more difficult to quantify, we also reiterate the band limitations of the X-ray facilities noted in Section I, which render them largely insensitive to photon energies $\gtrsim 8$ keV. Thus, both XMM-Newton and Chandra have little to no sensitivity to the hottest regions that may be present in the ICM, which could produce lower X-ray-derived temperatures for such regions.
V.2 Comparing rSZe and X-ray Measurements
In this merging system, we expect both a very hot on-average ICM temperature as well as spatial variations in the temperature. To simultaneously assess such a temperature structure, we separately fit the rSZe and X-ray data to a model that includes an overall mean temperature along with intrinsic scatter about this mean at the seven locations of the FTS pointings.